# Numerical optimization for the compatibility constant of the lasso

### v1

@@ Abstract
Compatibility condition and compatibility constant have been commonly used to evaluate the prediction error of the lasso when the number of variables exceeds the number of observations. However, the computation of the compatibility constant is generally difficult because it is a complicated nonlinear optimization problem. In this study, we present a numerical approach to compute the compatibility constant when the zero/nonzero pattern of true regression coefficients is given. We show that the optimization problem reduces to a quadratic program (QP) once the signs of the nonzero coefficients are specified. In this case, the compatibility constant can be obtained by solving QPs for all possible sign combinations. We also formulate a mixed-integer quadratic programming (MIQP) approach that can be applied when the number of true nonzero coefficients is moderately large. We investigate the finite-sample behavior of the compatibility constant for simulated data under a wide variety of parameter settings and compare the mean squared error with its theoretical error bound based on the compatibility constant. The behavior of the compatibility constant in finite samples is also investigated through a real data analysis.
Keywords: Lasso, Compatibility condition, Sparse estimation, Optimization
@@ 1 Introduction
The success of the lasso (least absolute shrinkage and selection operator; tibshirani1996lasso) lies in the computational efficiency (efron2004least; friedman2010regularization; boyd2011distributed; boyd2004convex; Stellato2020OSQP) and strong theoretical foundations in high-dimensional settings (bickel2009simultaneous; van2009conditions; buhlmann2011statistics; zhang2010nearly; zhao2006model; wainwright2009sharp; meinshausen2009lasso). In particular, the theory of the lasso has been shown to provide good prediction accuracy even when the number of variables exceeds the number of observations in practice. For example, some strong conditions, such as the compatibility condition, restricted eigenvalue condition, and irrepresentable condition, are sufficient conditions for obtaining a desirable property, such as oracle inequality and model selection consistency (zhao2006model; bickel2009simultaneous; van2009conditions; buhlmann2011statistics). The theory of the lasso shares many similarities with that of compressive sensing, where the restricted isometry property has been shown to guarantee sparse recovery (candes2005decoding; donoho2006compressed).
@@
This paper focuses on the compatibility condition (buhlmann2011statistics), one of the crucial conditions to investigate the oracle inequality of the lasso estimator. The compatibility condition can be weaker than other assumptions, including the restricted eigenvalue, irrepresentable, and restricted isometry conditions (van2009conditions; bickel2009simultaneous; buhlmann2011statistics). Therefore, it is crucial to investigate whether the compatibility condition holds; if it does not, there is a possibility that the lasso can fail to achieve good prediction accuracy.
The compatibility condition provides an upper bound of the mean squared error (MSE) in a high-dimensional setting. This upper bound is highly dependent on the compatibility constant, say $\phi$. Therefore, the computation of the compatibility constant is essential in understanding the prediction error of the lasso. Indeed, as the compatibility constant becomes large, the theoretical upper bound decreases, resulting in improvement of prediction accuracy (buhlmann2011statistics). Nevertheless, little attention has been paid to evaluating the compatibility constant for a given design matrix. One of the main reasons would be that most theoretical research has focused on establishing upper bounds for evaluating the prediction error, and the compatibility constant is typically used only to provide a sufficient condition for guaranteeing such upper bounds.
@@
However, understanding the finite-sample behavior of the compatibility constant would be essential for deepening the theoretical understanding of the lasso. For example, computing the compatibility constant enables us to investigate how large it is for a given design matrix, how it behaves with different numbers of observations and different numbers of dimensions, or how far the upper bound deviates from the prediction error. Such empirical investigations would provide a valuable step toward bridging the gap between the lasso theory and its practical performance.
In practice, the computation of $\phi$ is generally impossible as it depends on the zero/nonzero pattern of true regression coefficients that cannot be known in advance. If we assume that we know them in advance, the compatibility constant can be computed, but it can still be challenging in practice. The computation of the compatibility constant is similar to finding the smallest eigenvalue of the sample covariance matrix, but can differ and may be larger than the smallest eigenvalue as the corresponding vector (although not exactly an eigenvector) must lie within a nonconvex set defined by indices of true nonzero coefficients (buhlmann2011statistics). Therefore, it would be essential to develop a numerical approach to evaluate the compatibility constant in practice.
@@
In this study, we present a numerical approach to compute the compatibility constant when the zero/nonzero pattern of true regression coefficients is known. We show that when the signs of the nonzero coefficients are provided, the computation reduces to a Quadratic Programming (QP). Using this property, we can obtain the compatibility constant by solving QP for all combinations of signs and selecting the minimum value. When the number of true nonzero coefficients is moderately large, it is difficult to obtain the compatibility constant with the above QP-based procedure due to a huge number of combinations of the signs of nonzero regression coefficients. To handle the combinatorial complexity, we formulate the problem as a mixed-integer quadratic programming (MIQP, BertsimasTsitsiklis1997; BertsimasWeismantel2005) incorporating Big-$M$ constraints (NemhauserWolsey1988) or special ordered set type 1 (SOS1) variables (BealeTomlin1970; Williams2013).
@@
We investigate the numerical behavior of the compatibility constant for simulated data for a wide variety of parameter settings. The results show that the compatibility constant is highly dependent on the number of observations; therefore, we must take it into account when investigating the prediction error. Furthermore, we provide a heuristic approach to estimate the compatibility constant when the zero/nonzero pattern of true regression coefficients is unknown. This approach is applied to the analysis of the S&P500 dataset to demonstrate the relevance to the simulation results.
@@
The remainder of this paper is organized as follows. In Section 2, we briefly review the compatibility condition and its related topics. In Section 3, we provide a method for computing the compatibility constant based on QP and MIQP. Section 4 presents numerical simulations to compute the compatibility constant in a wide variety of settings, and compares the MSE and its upper bound constructed with the compatibility constant. In Section 5, we provide an actual data analysis to illustrate the behavior of the estimate of the compatibility constant. Section 6 concludes our work.
@@ 2 Preliminaries
where $\bm{\beta}$ is a regression coefficient vector and $\bm{\varepsilon}$ is an error vector. Consider the lasso estimation,
@@
For constructing the theory of the lasso for high-dimensional data, we assume that the true regression coefficients are sufficiently sparse. Let $S:=\{j\mid\beta_{j}\neq 0\}$, and let $s:=|S|$ denote its cardinality, which is assumed to be sufficiently small compared with $n$ and $p$. We may also consider the case $n<p$, but this is not necessary; we assume that $\frac{\log p}{n}\rightarrow 0$ as $n\rightarrow\infty$.
The compatibility condition is one of the most important assumptions for deriving an upper bound on the mean squared error (MSE) in high-dimensional settings. Let $\hat{\Sigma}=X^{\top}X/n$ be the Gram matrix of the design matrix. The compatibility condition, a condition for the design matrix $X$, is defined as follows:
@@
The compatibility condition (buhlmann2011statistics, p. 106) is met if there exists some $\phi>0$ such that, for any $\bm{v}\in\mathbb{R}^{p}$ satisfying $\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}$, it holds that
@@
The compatibility constant is related to the smallest eigenvalue of $\hat{\bm{\Sigma}}$ because $\|\bm{v}_{S}\|_{1}^{2}\leq s\|\bm{v}_{S}\|_{2}^{2}$. However, $\phi$ is generally larger than the smallest eigenvalue because of the constraint $\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}$ (buhlmann2011statistics). The constraint $\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}$ restricts the vectors to be in a nonconvex set where the main direction of the vector mostly aligns with $S$. In other words, feasible directions cannot deviate too far from the coordinates indexed by $S$.
To illustrate this intuition, consider the compound-symmetry structure, $\hat{\Sigma}=\rho\bm{1}_{p}\bm{1}_{p}^{\top}+(1-\rho)\bm{I}_{p}$ with $\rho$ being the correlation coefficient between any two predictors. When $\rho>0$, the largest eigenvalue of $\hat{\Sigma}$ is $1+(p-1)\rho$ and the corresponding eigenvector is $\bm{1}_{p}$. The other eigenvalues are all $1-\rho$ and the corresponding eigenvectors are those orthogonal to $\bm{1}_{p}$. However, a direction $\bm{1}_{p}$ cannot satisfy the constraint $\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}$ when $|S^{c}|\gg|S|$. Consequently, the feasible region excludes the direction associated with the largest eigenvalue, and any feasible $\bm{v}$ must deviate substantially from the direction $\bm{1}_{p}$. Therefore, when $\rho\approx 1$, $\bm{v}^{\top}\hat{\Sigma}\bm{v}$ that satisfies $\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}$ must be small, leading to a small compatibility constant. In contrast, when $\rho\approx 0$, the quadratic form $\bm{v}^{\top}\hat{\Sigma}\bm{v}$ does not vary significantly across different directions $\bm{v}$. As a result, the compatibility constant tends to be larger when the predictors are weakly correlated. In other words, the compatibility condition may hold when the predictors are not highly correlated.
Under the compatibility condition, we obtain the following result (Tibshirani2017Sparsity).
@@
For details of the result and its proof, please refer to Section 4.2 of Tibshirani2017Sparsity. We may obtain results whose constant value is slightly different from (3) (e.g., 9 might be changed to 4; buhlmann2011statistics), but the rate of convergence remains unchanged; roughly speaking, the error converges to 0 with high probability when $\log p/n\rightarrow 0$ as $n\rightarrow\infty$.
The compatibility constant $\phi$ plays an important role in evaluating the error bound. Since the upper bound in Proposition 4 is proportional to $\phi^{-2}$, the magnitude of the compatibility constant $\phi$ directly determines how large the bound of the MSE is. When $\phi$ is large, the upper bound becomes small, which implies that the lasso may achieve a meaningful prediction accuracy if the tuning parameter is properly chosen.
Several conditions other than the compatibility condition have been proposed for establishing theoretical properties of the lasso (van2009conditions). In particular, the irrepresentable condition and restricted eigenvalue condition are two popular conditions. Their purposes differ depending on the aspect of the problem. The irrepresentable condition is closely related to model selection consistency (zhao2006model), while the restricted eigenvalue condition and the restricted isometry property are often used to analyze prediction error (bickel2009simultaneous; candes2005decoding). On the other hand, the compatibility condition is related to oracle inequalities as described in Proposition 4. Among these conditions, the compatibility condition is known to be the weakest condition that still guarantees oracle inequalities. In fact, van2009conditions show that the compatibility condition is implied by both the restricted eigenvalue condition and the restricted isometry property. For this reason, the compatibility condition can be considered a flexible assumption for studying the prediction performance of the lasso in high-dimensional settings.
3 Numerical optimization for compatibility condition
@@
$\phi^{2}=\min_{\bm{v}}\left\{\frac{s\bm{v}^{\top}\hat{\Sigma}\bm{v}}{\|\bm{v}_{S}\|_{1}^{2}}\ \middle|\;\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}\right\}.$ (5)
The above problem turns out to be scale-invariant; define $\bm{v}^{\prime}=a\bm{v}$ with any nonzero value $a$. Then, we observe that $\bm{v}^{\prime}$ and $\bm{v}$ provide the same compatibility constant. Therefore, without loss of generality, we may restrict attention to vectors satisfying $\|\bm{v}_{S}\|_{1}=1$, in which case the constraint $\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}$ reduces to $\|\bm{v}_{S^{c}}\|_{1}\leq 3$. Hence, the problem (5) can be reformulated as
@@ 3.1 Quadratic Programming
is a convex set, we can easily obtain $\phi^{2}$ for fixed $\bm{z}$, say $\phi^{2}_{\bm{z}}$, by solving the following problem:
@@
All constraints in (8) are linear and the feasible region is convex. The objective function $\bm{v}^{\top}\hat{\Sigma}\bm{v}$ is also convex because $\hat{\Sigma}$ is positive semidefinite. Hence, for a fixed sign pattern, the optimization problem (8) becomes a standard convex quadratic program (QP). We can obtain $\phi^{2}_{\bm{z}}$ with a solver for quadratic programming, such as the OSQP library (Stellato2020OSQP). The compatibility constant is obtained by solving (8) for all possible sign patterns of $\bm{z}$ as below:

### v2

@@ Abstract
The compatibility constant plays an important role in evaluating the prediction error of the lasso in high-dimensional settings. However, the computation of the compatibility constant is generally difficult because it is a complicated nonconvex optimization problem. In this study, we present a numerical approach to compute the compatibility constant when the support of true regression coefficients is given. We show that the optimization problem reduces to a quadratic programming (QP) once the signs of the nonzero coefficients are specified. In this case, the compatibility constant can be obtained by solving QPs for all possible sign combinations. We also formulate a mixed-integer QP (MIQP) approach that can be applied when the number of true nonzero coefficients is large. We investigate the finite-sample behavior of the compatibility constant for simulated data under various parameter settings and compare the prediction error with its theoretical upper bound. The behavior of the compatibility constant in finite samples is also investigated through a real data analysis.
Keywords: Lasso, Compatibility constant, Sparse estimation, Quadratic Programming
@@ 1 Introduction
The success of the lasso (least absolute shrinkage and selection operator; Tibshirani, 1996) lies in the computational efficiency (Efron et al., 2004; Friedman et al., 2010; Boyd et al., 2011; Boyd and Vandenberghe, 2004; Stellato et al., 2020) and strong theoretical foundations in high-dimensional settings (Bickel et al., 2009; van de Geer and Bühlmann, 2009; Bühlmann and van de Geer, 2011; Zhang, 2010; Zhao and Yu, 2006; Wainwright, 2009; Meinshausen and Yu, 2009). In particular, the theory of the lasso has been shown to provide good estimation and prediction accuracy even when the number of variables exceeds the number of observations. For example, some regularity conditions, such as the compatibility condition, restricted eigenvalue condition, and irrepresentable condition, are sufficient conditions for obtaining a desirable property, such as oracle inequality and model selection consistency (Zhao and Yu, 2006; Bickel et al., 2009; van de Geer and Bühlmann, 2009; Bühlmann and van de Geer, 2011). The theory of the lasso shares many similarities with that of compressive sensing, where the restricted isometry property has been shown to guarantee sparse recovery (Candès and Tao, 2005; Donoho, 2006).
@@
This paper focuses on the compatibility condition (Bühlmann and van de Geer, 2011), one of the crucial conditions for establishing the oracle inequality of the lasso estimator. The compatibility condition is weaker than many other assumptions, including the restricted eigenvalue, irrepresentable, and restricted isometry conditions (van de Geer and Bühlmann, 2009; Bickel et al., 2009; Bühlmann and van de Geer, 2011). Therefore, it is crucial to investigate whether the compatibility condition holds; if it does not, there is a possibility that the lasso can fail to achieve good prediction accuracy.
@@
The compatibility condition provides an upper bound on prediction error in a high-dimensional setting. This upper bound is highly dependent on a quantity referred to as the compatibility constant. As the compatibility constant increases, the theoretical upper bound on the prediction error decreases (Bühlmann and van de Geer, 2011). Therefore, the computation of the compatibility constant is essential for understanding the finite-sample behavior of the upper bound on the prediction error. Once we know how to compute the compatibility constant, we can evaluate not only the upper bound on the prediction error but also how far the upper bound deviates from the actual prediction error. In addition, we can study how the compatibility constant changes with the number of observations and the dimensionality. Such empirical investigations would provide a valuable step toward bridging the gap between lasso theory and its practical performance.
Nevertheless, the compatibility constant has rarely been computed for a given design matrix. This is partly because most theoretical research has focused on establishing upper bounds for evaluating the prediction error, and the compatibility constant is typically used only to provide a sufficient condition for guaranteeing such upper bounds. In other words, the compatibility constant is often treated as a theoretical quantity rather than being evaluated numerically for specific design matrices. Another reason is the intrinsic difficulty in computing the compatibility constant. First, the computation of the compatibility constant is generally impossible, as it depends on the support of the true regression coefficients (i.e., the set of indices corresponding to nonzero coefficients), which cannot be known in advance. Even if we know the support in advance, the computation is still challenging in practice. The computation is similar to finding the smallest eigenvalue of the sample covariance matrix but can differ and may be significantly larger than this eigenvalue. This is because the associated vector (though not exactly an eigenvector) must lie within a nonconvex set defined by the support of the true nonzero coefficients (Bühlmann and van de Geer, 2011). As a result, an analytical approach is generally infeasible. Although a numerical approach is helpful, the problem reduces to a complicated nonconvex optimization problem.
@@
In this study, we present a numerical approach to compute the compatibility constant when the support of true regression coefficients is known. We show that when the signs of the nonzero coefficients are provided, the computation reduces to a Quadratic Programming (QP). Using this property, the compatibility constant can be obtained by solving QPs for all possible sign combinations and selecting the smallest value. When the number of true nonzero coefficients is large, it is difficult to obtain the compatibility constant with the above QP-based procedure due to a huge number of combinations of the signs of the true nonzero regression coefficients. To handle the combinatorial complexity, we formulate the problem as a mixed-integer quadratic programming (MIQP, Bertsimas and Tsitsiklis, 1997; Bertsimas and Weismantel, 2005) incorporating Big-$M$ constraints (Nemhauser and Wolsey, 1988) or special ordered set type 1 (SOS1) variables (Beale and Tomlin, 1970; Williams, 2013).
@@
We investigate the numerical behavior of the compatibility constant for simulated data under a wide variety of parameter settings. The results show that the compatibility constant is highly dependent on the number of observations; therefore, we must take it into account when investigating the prediction error. We also investigate how the bound on the prediction error changes at each step of the proof of the oracle inequality. Furthermore, we provide a heuristic approach to estimate the compatibility constant when the support of true regression coefficients is unknown. This approach is applied to the analysis of the S&P500 dataset to demonstrate the relevance to the simulation results.
@@
The remainder of this paper is organized as follows. In Section 2, we briefly review the compatibility condition and its related topics. In Section 3, we provide a method for computing the compatibility constant based on QP and MIQP. Section 4 presents numerical simulations to compute the compatibility constant under a wide variety of settings, and compares the prediction error and its upper bound constructed with the compatibility constant. In Section 5, we provide an actual data analysis to illustrate the behavior of the estimate of the compatibility constant. Section 6 concludes our work.
@@ 2 Preliminaries
where $\bm{\beta}$ is a regression coefficient vector and $\bm{\varepsilon}$ is an error vector. Consider the lasso problem,
@@
Let $S:=\{j\mid\beta_{j}\neq 0\}$, and let $s:=|S|$ denote its cardinality. For constructing the theory of the lasso for high-dimensional data, we assume that $s$ is sufficiently small compared with $n$ and $p$; in other words, true regression coefficients are sufficiently sparse. We may also consider the case $n<p$, but this is not necessary; we assume that $\frac{\log p}{n}\rightarrow 0$ as $n\rightarrow\infty$.
The compatibility condition is one of the most important assumptions for deriving an upper bound on the prediction error in high-dimensional settings. Let $\hat{\Sigma}=X^{\top}X/n$ be the Gram matrix of the design matrix. The compatibility condition, a condition for the design matrix $X$, is defined as follows:
@@
The compatibility condition (Bühlmann and van de Geer, 2011, p. 106) is met if there exists some $\phi>0$ such that, for any $\bm{v}\in\mathbb{R}^{p}$ satisfying $\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}$, it holds that
@@
The compatibility constant is related to the smallest eigenvalue of $\hat{\Sigma}$ because $\|\bm{v}_{S}\|_{1}^{2}\leq s\|\bm{v}_{S}\|_{2}^{2}$. However, $\phi$ is generally larger than this eigenvalue because of the constraint $\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}$ (Bühlmann and van de Geer, 2011). The constraint $\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}$ restricts the vectors to be in a nonconvex set where the main direction of the vector mostly aligns with $S$. In other words, feasible directions cannot deviate too far from the coordinates indexed by $S$.
To illustrate this intuition, consider the compound-symmetry structure, $\hat{\Sigma}=\rho\bm{1}_{p}\bm{1}_{p}^{\top}+(1-\rho)I_{p}$ with $\rho$ being the correlation coefficient between any two predictors. When $\rho>0$, the largest eigenvalue of $\hat{\Sigma}$ is $1+(p-1)\rho$ and the corresponding eigenvector is $\bm{1}_{p}$. The other eigenvalues are all $1-\rho$ and the corresponding eigenvectors are those orthogonal to $\bm{1}_{p}$. However, a direction $\bm{1}_{p}$ cannot satisfy the constraint $\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}$ when $|S^{c}|\gg|S|$. Consequently, the feasible region excludes the direction associated with the largest eigenvalue, and any feasible $\bm{v}$ must deviate substantially from the direction $\bm{1}_{p}$. Therefore, when $\rho\approx 1$, $\bm{v}^{\top}\hat{\Sigma}\bm{v}$ that satisfies $\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}$ must be small, leading to a small compatibility constant. In contrast, when $\rho\approx 0$, the quadratic form $\bm{v}^{\top}\hat{\Sigma}\bm{v}$ does not vary significantly depending on the direction of $\bm{v}$. As a result, the compatibility constant tends to be larger.
Under the compatibility condition, we obtain the following result (Tibshirani and Wasserman, 2017).
@@
For details of the result and its proof, please refer to Section 4.2 of Tibshirani and Wasserman (2017). Some of the proofs are also given in Appendix B. We may obtain results whose constant value is slightly different from (3) (e.g., 9 might be changed to 4; Bühlmann and van de Geer, 2011), but the rate of convergence remains unchanged; roughly speaking, the error converges to 0 with high probability when $\log p/n\rightarrow 0$ as $n\rightarrow\infty$.
The compatibility constant $\phi$ plays an important role in evaluating the error bound. Since the upper bound in Proposition 4 is proportional to $\phi^{-2}$, the magnitude of the compatibility constant $\phi$ directly determines how large the bound of the prediction error is. When $\phi$ is large, the upper bound becomes small, which implies that the lasso may achieve a meaningful prediction accuracy if the tuning parameter is properly chosen.
Several conditions other than the compatibility condition have been proposed for establishing theoretical properties of the lasso (van de Geer and Bühlmann, 2009). In particular, the irrepresentable condition and restricted eigenvalue condition are two popular conditions. Their purposes differ depending on the aspect of the problem. The irrepresentable condition is closely related to model selection consistency (Zhao and Yu, 2006), while the restricted eigenvalue condition and the restricted isometry property are often used to analyze prediction error (Bickel et al., 2009; Candès and Tao, 2005). On the other hand, the compatibility condition is related to oracle inequalities as described in Proposition 4. Among these conditions, the compatibility condition is known to be the weakest condition that still guarantees oracle inequalities. In fact, van de Geer and Bühlmann (2009) show that the compatibility condition is implied by both the restricted eigenvalue condition and the restricted isometry property. For this reason, the compatibility condition can be considered a flexible assumption for studying the prediction performance of the lasso in high-dimensional settings.
3 Numerical optimization for compatibility constant
@@
$\phi^{2}=\min_{\{\bm{v}\mid\bm{v}\neq\bm{0}\}}\left\{\frac{s\bm{v}^{\top}\hat{\Sigma}\bm{v}}{\|\bm{v}_{S}\|_{1}^{2}}\ \middle|\;\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}\right\}.$ (5)
The above problem turns out to be scale-invariant. Indeed, for any nonzero scalar $a$, the vectors $\bm{v}$ and $a\bm{v}$ yield the same compatibility constant. Therefore, without loss of generality, we may impose the constraint $\|\bm{v}_{S}\|_{1}=1$, in which case the constraint $\|\bm{v}_{S^{c}}\|_{1}\leq 3\|\bm{v}_{S}\|_{1}$ reduces to $\|\bm{v}_{S^{c}}\|_{1}\leq 3$. Hence, the problem (5) can be reformulated as
@@ 3.1 Quadratic Programming
is a convex set, we can easily obtain $\phi^{2}$ for given $\bm{z}$, say $\phi^{2}_{\bm{z}}$, by solving the following problem:
@@
All constraints in (7) are linear and the feasible region is convex. The objective function $\bm{v}^{\top}\hat{\Sigma}\bm{v}$ is also convex because $\hat{\Sigma}$ is positive semidefinite. Hence, for a fixed sign pattern, the optimization problem (7) becomes a standard convex quadratic program (QP). We can obtain $\phi^{2}_{\bm{z}}$ with a solver for quadratic programming, such as the OSQP library (Stellato et al., 2020). The compatibility constant is obtained by solving (7) for all possible sign patterns of $\bm{z}$ as below: